How long will it take workers A and B, together, to finish a job which can be done by A alone in 6 days, by B alone in 4 days?: How long will it take workers A and B, together, to finish a job which can be done by A alone in 6 days, by B alone in 4 days?
Answer by edjones(2391)

a does 1/6 of the job in 1 day.
b does 1/4 of the job in 1 day.
.
1/6 + 1/4 =
2/12 + 3/12 =
5/12 of the job in 1 day working together.
So, they can complete the job in 12/5 days or 2 2/5 days.
.
Ed
Question 152195: How long will it take workers A and B, together, to finish a job which can be done by A alone in 6 days, by B alone in 4 days?: How long will it take workers A and B, together, to finish a job which can be done by A alone in 6 days, by B alone in 4 days?
Answer by orca(336)

This is a shared work problem.
Let the total work be 1.
Worker A can do 1/6 of the total work in one day.
Worker B can do 1/4 of the total work in one day.
Working jointly, they can do 1/6 + 1/4 = 2/12 + 3/12 = 5/12 of the total work.
So it takes them 1/(5/12) = 12/5 days to finish the job.
Question 152195: How long will it take workers A and B, together, to finish a job which can be done by A alone in 6 days, by B alone in 4 days?: How long will it take workers A and B, together, to finish a job which can be done by A alone in 6 days, by B alone in 4 days?
Answer by mducky2(55)

Since it took 6 days for A to finish a job and 4 days for B to finish a job, we can represent the rate of each worker as:
A = 1/6 job per day
B = 1/4 job per day

We can represent the problem as:
Number of days * (Rate of worker A + Rate of worker B) = One job

[x(1/6 + 1/4)] = 1
[x(2/12 + 3/12)] = 1
x(5/12) = 1
x = 12/5
x = 2.4 days